OpenStudy (anonymous):

Which of the following is the equation of a line passes through (-2,-1) and (-4,-3)? a. y=1/2 x+1 b.y=x+1 i pick this one c.y=1/2 x-1 d.y=x-1

OpenStudy (anonymous):

b is correct

OpenStudy (anonymous):

how do u know

OpenStudy (anonymous):

the coordinates are in the form (x,y), so plug in the values y=x+1 -1=-2+1 -1=-1

OpenStudy (anonymous):

slope= change in y over change in x

myininaya (myininaya):

\[\frac{-2-(-4)}{-1-(-3)}=\frac{2}{2}=1\] y=mx+b y=1x+b y=x+b -1=-2+b -1+2=b 1=b y=x+1

OpenStudy (anonymous):

\[slope=\frac{y1-y2}{x1-x2}\]

OpenStudy (anonymous):

Just follow Myininaya from there :)

OpenStudy (anonymous):

u guys are smart damn i need tutoring from yall lolol

OpenStudy (anonymous):

u guys are smart damn i need tutoring from yall lolol